OpenStudy (anonymous):

Rewrite the equation in slope=intercept form. 6x-3y=-9 .

OpenStudy (anonymous):

I think the most useful form of straight-line equations is the "slope-intercept" form: y = mx + b In order to get the equation, 6x - 3y = -9 into that form, you'd have to get y on one side, right? Okay, so in order to do that, you need to move the other integers to the other side of the equal sign. 6x-3y = -9 Let's first subtract the 3. 6x-y = -9 - 3 (= - 12) Then let's subtract the 6x. y = -12 - 6x. Does that make sense and answer your question?

OpenStudy (anonymous):

Yes it does . Can I have help on one more ?

OpenStudy (anonymous):

Yeah, I'll try to help. What's the problem?

OpenStudy (anonymous):

Sorry I'm just replying back. I didn't know that you replied back but anyway . 2x+5y=-10

OpenStudy (anonymous):

Okay, so you want to get it into the form y = mx + b So in this problem, you'd need to get "y" alone on one side of the equal sign. 2x + 5y = -10 First, you'd want to subtract 2x from both sides. 2x (-2x) + 5y = - 10 - 2x You always want to put the variable first so, you'd switch the variables around to make.. 5y = -2x - 10 Alright, and then you need to divide 5 from both sides. 5y (/5) = -2/5x - 10/5 Then it'd be y = -2/5x - 2 Sorry, I typed that from the top of my head; I hope that made sense. I'll clarify if it doesn't.